# LCES: Zero-shot Automated Essay Scoring via Pairwise Comparisons Using Large Language Models

### v1

@@ 1 Introduction
To address the issues above, we explore an alternative AES formulation. Instead of predicting absolute scores, we instruct the LLM to perform pairwise comparisons in which the better of two essays is determined. This approach is inspired by recent advances in LLM-based evaluation for natural language generation (Liu et al., 2024; Liusie et al., 2024a), dialogue systems (Park et al., 2024), and information retrieval (Qin et al., 2024), where pairwise comparisons have demonstrated stronger alignment with human preferences. Despite its promise, pairwise comparisons remain largely unexplored in the AES literature.
@@
The contributions of this work are summarized as follows: (1) We introduce the first AES framework based on LLM-generated pairwise comparisons, addressing key limitations of direct score generation. (2) We leverage RankNet to convert LLM-generated preferences into continuous scores, enabling accurate and computationally efficient zero-shot AES. (3) Extensive experiments confirm that LCES outperforms conventional zero-shot AES baselines and is robust across different types of LLMs.
@@ 3.1 Pairwise Comparison Generation
To apply this comparison procedure, we construct a set of essay pairs. Let =$\mathcal{I}=\{(i,j)\mid i\neq j,\ i,j\in\{1,2,\ldots,N\}\}$ be the set of all possible ordered essay pairs. Since comparing all $N(N-1)$ pairs is computationally expensive, we randomly sample a subset $\mathcal{I}_{s}\subset\mathcal{I}$ containing $M$ pairs, where $M\ll N(N-1)$. For each sampled pair, we obtain a debiased label $\tilde{c}_{ij}$ as described above. This yields the pairwise comparison dataset $\mathcal{D}_{\text{pair}}=\{(x_{i},x_{j},\tilde{c}_{ij})\mid(i,j)\in\mathcal{I}_{s}\}$, which is used to train the RankNet model.
@@ 3.3 Output Conversion
To produce a score $\hat{y}_{i}$ within a rubric-defined range $[y_{\min},y_{\max}]$, we apply a linear transformation to the latent scores:
@@ 4.3 Experimental Setup
Implementation Details. The number of sampled pairwise comparisons $M$ was set to 5,000 to construct the $\mathcal{D}_{\text{pair}}$ dataset. Essay embedding vectors $\bm{h}_{i}$ were generated using OpenAI’s text-embedding-v3-large model. Results obtained with alternative embedding models are presented in Appendix B. The RankNet model was trained for 100 epochs using the Adam Kingma and Ba (2015) optimizer with a learning rate of 0.001. The full set of hyperparameters is provided in Appendix C.
@@ 4.4 Results and Discussion
Moreover, LCES demonstrates strong robustness across backbone LLMs. On ASAP, the standard deviation of its average performance across five backbone models is just 0.021, compared to 0.072 for MTS and 0.122 for Vanilla. On TOEFL11, LCES similarly shows low variability, with a standard deviation of 0.048 across models, outperforming MTS (0.106) and Vanilla (0.079). These low inter-model variances indicate that LCES remains stable regardless of backbone choice, whereas MTS and Vanilla fluctuate more, making their performance less predictable.
In summary, LCES exhibited high zero-shot accuracy across diverse settings, with robustness to LLM choice and reliable scaling with model size. Even when the model struggles to distinguish essay quality, as with Mistral-7B on TOEFL11, LCES remains competitive, supporting its practicality in real-world AES scenarios.
@@ 5.1 Comparison with Supervised Models
As shown in Table 3, LCES with Llama-3.1-8B, which achieved the highest overall performance among all tested LLMs in the zero-shot experiments (see Section 4.4), obtains QWK scores that are comparable to several supervised learning models. While NPCR, BERT, and RoBERTa still outperform LCES, the performance gap has significantly narrowed in comparison with previously reported zero-shot methods. In addition, LCES achieves performance on par with the strong cross-prompt baselines PAES and PMAES [^3] PMAES was run with a smaller batch size due to GPU limitations (RTX 4090), which may have led to reduced performance.. This level of performance is unprecedented among zero-shot AES methods. These results highlight the effectiveness of the proposed method in the absence of scored essays.
@@ 5.4 Performance in the Inductive Setting
In LCES, pairwise preferences predicted by an LLM are used to train a scoring function, $f$. When all target essays are available at once, comparisons can be directly made among them. We refer to this as the transductive setting, which corresponds to the main experimental setup used throughout this paper. In contrast, the inductive setting assumes that new essays must be scored individually, without additional comparisons. Instead, the learned function $f$, trained on prior comparisons, is applied for score estimation. Because $f$ maps essay embeddings to scalar scores, it can generalize to new essays without further pairwise information.
To simulate this scenario, we train the scoring function $f$ on pairwise comparisons constructed from 90% of the essays in each dataset (ASAP or TOEFL11) and use it to predict scores for the remaining 10%. We use GPT-4o-mini for its computational efficiency and low API cost.
@@ 6 Conclusion
Despite its advantages, LCES has several limitations. First, it relies on pairwise preference labels generated by an LLM, which may contain noise or inconsistencies. These imperfect labels directly affect the quality of learned scoring function $f$.
@@
Finally, the zero-shot nature of LCES means that, without labeled data, its performance cannot be quantitatively assessed. For practical deployment, this necessitates manually scoring a subset of essays to establish a benchmark for evaluation.
@@ Acknowledgments
This research was conducted as part of an internal project at Deloitte Touche Tohmatsu LLC. The authors would like to thank Tomoaki Geka and Tomotake Kozu for helpful discussions and support.
@@ A.1 ASAP
User Prompt # Prompt {prompt} # Rubric Guidelines {rubric} # Note I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER).
The relevant entities are identified in the text and then replaced with a string such as "PERSON", "ORGANIZATION", "LOCATION", "DATE", "TIME", "MONEY", "PERCENT”, “CAPS” (any capitalized word) and “NUM” (any digits). Please do not penalize the essay because of the anonymizations. # Essay1 {essay1} # Essay2 {essay2} Provide your reasoning and final decision in json format: { "reasoning": "Your reasoning in one sentence here.", "preference": "essay1" or "essay2" or "tie" }
@@ A.2 TOEFL11
User Prompt # Prompt {prompt} # Rubric Guidelines {rubric} # Essay1 {essay1}
# Essay2 {essay2} Provide your reasoning and final decision in json format: { "reasoning": "Your reasoning in one sentence here.", "preference": "essay1" or "essay2" or "tie" }
@@ Appendix B Embedding Models
text-embedding-v3-large 0.653 0.645
text-embedding-v3-small 0.668 0.630
@@
We compare four pretrained embedding models used to convert essays into fixed-length vectors for RankNet. Two of them are OpenAI models: text-embedding-v3-large (3072 dimensions) and text-embedding-v3-small (1536 dimensions), both of which were designed for semantic similarity tasks. The other two are BERT-base and RoBERTa-base. For these models, we use the [CLS] token from the final hidden layer as the essay representation.

### v2

@@ 1 Introduction
Since such problems exist with direct scoring, this study explores an alternative evaluation method, namely pairwise evaluation. Instead of predicting absolute scores, we instruct the LLM to perform pairwise comparisons in which the better of two essays is determined. This approach is inspired by recent advances in LLM-based evaluation for natural language generation (Liu et al., 2024; Liusie et al., 2024a), dialogue systems (Park et al., 2024), and information retrieval (Qin et al., 2024), where pairwise comparisons have demonstrated stronger alignment with human preferences. Despite its promise, pairwise comparisons remain largely unexplored in the AES literature.
@@
The contributions of this work are summarized as follows:
  1. (1)
    We introduce the first AES framework based on LLM-generated pairwise comparisons, addressing key limitations of direct score generation.
  2. (2)
    We leverage RankNet to convert LLM-generated preferences into continuous scores, enabling accurate and computationally efficient zero-shot AES.
  3. (3)
    Extensive experiments confirm that LCES outperforms conventional zero-shot AES baselines and is robust across different types of LLMs.
@@ 3.1 Pairwise Comparison Generation
To apply this comparison procedure, we construct a set of essay pairs. Let $\mathcal{I}=\{(i,j)\mid i\neq j,\ i,j\in\{1,2,\ldots,N\}\}$ be the set of all possible ordered essay pairs. Since comparing all N(N1)N(N-1) pairs is computationally expensive, we randomly sample a subset $\mathcal{I}_{s}\subset\mathcal{I}$ containing $M$ pairs, where $M\ll N(N-1)$. For each sampled pair, we obtain a debiased label $\tilde{c}_{ij}$ as described above. This yields the pairwise comparison dataset $\mathcal{D}_{\text{pair}}=\{(x_{i},x_{j},\tilde{c}_{ij})\mid(i,j)\in\mathcal{I}_{s}\}$, which is used to train the RankNet model.
@@ 3.3 Output Conversion
To produce a score $\hat{y}_{i}$ within a rubric-defined range $y_{\min},y_{\max}]$, we apply a linear transformation to the latent scores:
@@ 4.3 Experimental Setup
Implementation Details. The number of sampled pairwise comparisons $M$ was set to 5,000 to construct the $\mathcal{D}_{\text{pair}}$ dataset. Essay embedding vectors $\bm{h}_{i}$ were generated using OpenAI’s text-embedding-3-large model. Results obtained with alternative embedding models are presented in Appendix B. The RankNet model was trained for 100 epochs using the Adam Kingma and Ba (2015) optimizer with a learning rate of 0.001. The full set of hyperparameters is provided in Appendix C.
@@ 4.4 Results and Discussion
Moreover, LCES exhibited lower performance variance across different LLM backbones compared to conventional methods. On ASAP, the standard deviation of its average performance across five backbone models is just 0.021, compared to 0.072 for MTS and 0.122 for Vanilla. On TOEFL11, LCES similarly shows low variability, with a standard deviation of 0.048 across models, outperforming MTS (0.106) and Vanilla (0.079). These low inter-model variances indicate that LCES remains stable regardless of backbone choice, whereas MTS and Vanilla fluctuate more, making their performance less predictable.
@@ 5.1 Comparison with Supervised Models
As shown in Table 3, LCES with Llama-3.1-8B, which achieved the highest overall performance among all tested LLMs in the zero-shot experiments (see Section 4.4), obtains QWK scores that are comparable to several supervised learning models. While NPCR, BERT, and RoBERTa still outperform LCES, the performance gap has significantly narrowed in comparison with previously reported zero-shot methods. In addition, LCES with Llama-3.1-8B achieves performance on par with the strong cross-prompt baselines PAES and PMAES [^3] PMAES was run with a smaller batch size due to GPU limitations (RTX 4090), which may have led to reduced performance.. Indeed, a Wilcoxon signed-rank test revealed no statistically significant differences in QWK scores between LCES and either PAES or PMAES ($p$-values all above the 0.05 significance threshold). This level of performance is unprecedented among zero-shot AES methods. These results highlight the effectiveness of the proposed method in the absence of scored essays.
@@ 5.4 Performance in the Inductive Setting
In this section, we evaluate the performance of LCES under different problem settings. Throughout this paper, we have considered a problem setting where all target essays are available at once for scoring–we refer to this as the transductive setting. However, another scenario is possible where new essays arrive individually for scoring after the model has been trained–we refer to this as the inductive setting. The key advantage of LCES in the inductive setting is that the scoring model $f$, learned during RankNet training, maps essay embeddings to scalar scores and can thus generalize to unseen essays without requiring additional pairwise comparisons.
To simulate this scenario, we train the scoring model $f$ within the LCES framework on pairwise comparisons constructed from 90% of the essays in each dataset (ASAP or TOEFL11), and then use it to predict scores for the remaining 10%. We use GPT-4o-mini for its computational efficiency and low API cost.
@@ 6 Conclusion
Despite its advantages, LCES has several limitations. First, it relies on pairwise preference labels generated by an LLM, which may contain noise or inconsistencies. These imperfect labels directly affect the quality of learned scoring model $f$.
@@
Finally, the zero-shot nature of LCES means that, without labeled data, its performance cannot be quantitatively assessed. For practical deployment, caution is warranted when used in high-stakes exams.
@@ Acknowledgments
This research was conducted as part of an internal project at Deloitte Touche Tohmatsu LLC. The authors would like to thank Tomoaki Geka and Tomotake Kozu for helpful discussions and support. We also used ChatGPT to assist in clarifying the structure and wording of certain parts of the manuscript.
@@ A.1 ASAP
User Prompt # Prompt {prompt} # Rubric Guidelines {rubric} # Note I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as “PERSON", “ORGANIZATION", “LOCATION", “DATE", “TIME", “MONEY", “PERCENT”, “CAPS” (any capitalized word) and “NUM” (any digits). Please do not penalize the essay because of the anonymizations. # Essay1 {essay1} # Essay2 {essay2} Provide your reasoning and final decision in json format: { "reasoning": "Your reasoning in one sentence here.", "preference": "essay1" or "essay2" or "tie" }
@@ A.2 TOEFL11
User Prompt # Prompt {prompt} # Rubric Guidelines {rubric} # Essay1 {essay1} # Essay2 {essay2} Provide your reasoning and final decision in json format: { "reasoning": "Your reasoning in one sentence here.", "preference": "essay1" or "essay2" or "tie" }
@@ Appendix B Embedding Models
We compare four pretrained embedding models used to convert essays into fixed-length vectors for RankNet. Two of them are OpenAI models: text-embedding-3-large (3072 dimensions) and text-embedding-3-small (1536 dimensions), both of which were designed for semantic similarity tasks. The other two are BERT-base and RoBERTa-base. For these models, we use the [CLS] token from the final hidden layer as the essay representation.
@@
text-embedding-3-large 0.653 0.645
text-embedding-3-small 0.668 0.630